Grade 11Mechanics

A cyclist riding a bicycle at a speed of 5.1 m/s round a curve at an angle of 53^o with the horizontal in order to avoid toppling over. Calculate the radius of the curve.

dee
7 Years agoGrade 11

1 Answer

Eshan
7 Years ago

To find the radius of the curve that a cyclist must navigate to avoid toppling over while riding at a speed of 5.1 m/s and at an angle of 53 degrees, we can use some principles of physics, particularly those related to circular motion and forces acting on the cyclist.

Understanding Forces in Circular Motion

When a cyclist rounds a curve, they experience a centripetal force that keeps them moving in a circular path. This force is provided by the friction between the tires and the road. Additionally, the angle of the cyclist's lean plays a crucial role in maintaining balance and preventing toppling.

Key Concepts

  • Centripetal Force (Fc): The force that keeps an object moving in a circular path.
  • Weight (W): The gravitational force acting on the cyclist (W = mg, where m is mass and g is acceleration due to gravity).
  • Banking Angle (θ): The angle at which the cyclist leans into the curve.
  • Radius (r): The radius of the curve.

Applying the Relevant Formula

The relationship between these variables can be understood through the equation for centripetal acceleration:

Fc = (mv^2)/r

Where:

  • Fc is the centripetal force,
  • m is the mass of the cyclist,
  • v is the speed (5.1 m/s),
  • r is the radius of the curve.

In this situation, the cyclist leans against gravitational force and centripetal force. The forces can be broken down into components:

Force Components

The vertical component of the gravitational force must equal the vertical component of the normal force, while the horizontal component provides the necessary centripetal force.

Using trigonometry:

  • The vertical force balance: W = N cos(θ)
  • The horizontal force (centripetal): Fc = N sin(θ)

Combining the Equations

From these equations, we can express the normal force (N) in terms of weight:

N = W / cos(θ)

Substituting this into the centripetal force equation gives:

Fc = (W / cos(θ)) sin(θ)

Replacing W with mg, we have:

Fc = (mg / cos(θ)) sin(θ)

Finding the Radius

Now we can set this equal to the centripetal force expression:

(mv^2) / r = (mg / cos(θ)) sin(θ)

Rearranging gives:

r = (v^2 * cos(θ)) / (g * sin(θ))

Calculating Values

Inserting the known values:

  • v = 5.1 m/s
  • g = 9.81 m/s² (acceleration due to gravity)
  • θ = 53 degrees (convert to radians if needed)

Using the trigonometric values:

  • cos(53°) ≈ 0.6018
  • sin(53°) ≈ 0.7986

Now plugging in these values:

r = (5.1² * 0.6018) / (9.81 * 0.7986)

Calculating:

r = (26.01 * 0.6018) / (7.8347)

r ≈ 2.01 m

Final Thoughts

The radius of the curve that the cyclist must navigate, while maintaining balance and avoiding toppling over, is approximately 2.01 meters. This calculation illustrates how physics principles, including forces and motion, interact in real-world scenarios like cycling around a curve.